Find the third degree polynomial function in standard form with leading term coefficient 2,and zeros -1, 2, 3.
Answer by ikleyn(52803)

.
f(x) = 2*(x-(-1))*(x-2)*(x-3) =  (or, which is the same) = 2*(x+1)*(x-2)*(x-3).


I leave it to you to open parentheses and to get the standard form.


Making calculations is a students' privilege and a honor duty.